# Transferring Features Across Language Models With Model Stitching

### v1

@@ 5 Feature Analysis
$\operatorname{Corr}_{t_{\tau}}\left(f_{i}\left(h_{A}^{\ell_{A}}(t_{\tau});\theta\right)v_{i,t_{\tau+1,A}},f_{i}\left(h^{\ell_{B}}_{B}(t_{\tau});\theta^{\prime}\right)v^{\prime}_{i,t_{\tau+1,B}}\right)$ (10)
@@ 7 Conclusion
We find that we can faithfully stitch between language models of different sizes of the same family, which can be leveraged to transfer SAEs, probes, and steering vectors. The SAEs can be trained to convergence faster than training from scratch, demonstrating an application that uses universality to benefit SAE training. Probes and steering vectors can also be transferred with no additional training in specific cases. Finally, we perform a case study on GPT2 and find differences between how the stitch transfers semantic and structural features and discover that specific functional features have their roles preserved.
@@ Appendix A Stitch Training Details
The stitch training methodology is quite minimal. For the training dataset, we collect activations at the desired layers in both models over the first $180$k samples of OpenWebText with a context size of $512$ ($128$ for Gemma due to compute constraints). We mask out all special tokens. The stitches themselves are two separate Linear layers (with biases initialized to 00) that map between the dimensions of $A$ and $B$. We use the Adam optimizer with a learning rate of 1e-4 and clip gradient norms to $1.0$. We found minimal sensitivity to learning rate schedule, so we just use a cosine annealing decay, and found that $2$ epochs is sufficient for convergence (though even $1$ is probably enough).
@@ B.1 SAE Methodology
The insight is that we can think of transferring a sparse autoencoder from model $A$ to model $B$ as capturing the following computation. For every latent in model $B$ $\ell_{B}\in\mathbb{R}^{d_{B}}$:
@@
    Stitch $\ell_{B}$ to $A$ using $\mathcal{T}_{\downarrow}$.
@@
    Stitch the reconstruction $\textbf{SAE}(x;\theta)$ back to $B$ using $\mathcal{T}_{\uparrow}$.
@@
$(\mathcal{T}_{\uparrow}\circ\textbf{SAE}\circ\mathcal{T}_{\downarrow})(\ell_{B})$ $=\textbf{SAE}(\ell_{B}P_{\downarrow}+b_{\downarrow};\theta)P_{\uparrow}+b_{\uparrow}$
$=\Big{(}\sigma([\ell_{B}P_{\downarrow}+b_{\downarrow}]W_{e}+b_{e})W_{d}+b_{d}\Big{)}P_{\uparrow}+b_{\uparrow}$
$=\sigma(\ell_{B}[P_{\downarrow}W_{e}]+[b_{\downarrow}W_{e}+b_{e}])[W_{d}P_{\uparrow}]+[b_{d}P_{\uparrow}+b_{\uparrow}]$
$=\textbf{SAE}(\ell_{B};\theta^{\prime}),$
@@
$\|\textbf{SAE}(\ell_{B};\theta^{\prime})-\ell_{B}\|_{2}$ (13)
@@ B.2 Probes and Steering Vectors
To transfer a linear probe from $A$ to $B$, we simply apply the probe on the down-stitched residual stream $\mathcal{T}_{\downarrow}(h_{B}^{\ell_{B}}(t))$. If the probe is linear and has normal direction $w$, it is equivalent to using the vector $wP_{\downarrow}^{T}$ in $B$. Steering vectors are just transferred directly using $P_{\uparrow}$ without the bias.

### v2

@@ 5 Feature Analysis
$\operatorname{Corr}_{t_{\tau}}\left(f_{i}\left(A^{\ell_{A}}(t_{\tau});\theta\right)v_{i,t_{\tau+1,A}},f_{i}\left(B^{\ell_{B}}(t_{\tau});\theta^{\prime}\right)v^{\prime}_{i,t_{\tau+1,B}}\right)$ (10)
@@ 7 Conclusion
We find that we can faithfully stitch between language models of different sizes of the same family, which can be leveraged to transfer SAEs, probes, and steering vectors. The SAEs can be trained to convergence faster than training from scratch, demonstrating an application that uses weak universality to benefit SAE training. Probes and steering vectors can also be transferred with no additional training in specific cases. Finally, we perform a case study on GPT2 and find differences between how the stitch transfers semantic and structural features and discover that specific functional features have their roles preserved.
@@ Appendix A Stitch Training Details
The stitch training methodology is quite minimal. For the training dataset, we collect activations at the desired layers in both models over the first $180$k samples of OpenWebText with a context size of $512$ ($128$ for Gemma due to compute constraints) and evaluate over the next $1$k samples with the same context size. We mask out all special tokens. The stitches themselves are two separate Linear layers (with biases initialized to 00) that map between the dimensions of $A$ and $B$. We use the Adam optimizer with a learning rate of 1e-4 and clip gradient norms to $1.0$. We found minimal sensitivity to learning rate schedule, so we just use a cosine annealing decay, and found that $2$ epochs is sufficient for convergence (though even $1$ is probably enough).
@@ B.1 SAE Methodology
The insight is that we can think of transferring a sparse autoencoder from model $A$ to model $B$ as capturing the following computation. For every latent in model $B$ $h_{B}\in\mathbb{R}^{d_{B}}$:
@@
    Stitch $h_{B}$ to $A$ using $\mathcal{T}_{\downarrow}$.
@@
    Stitch the reconstruction back to $B$ using $\mathcal{T}_{\uparrow}$.
@@
$(\mathcal{T}_{\uparrow}\circ\textbf{SAE}\circ\mathcal{T}_{\downarrow})(h_{B})$ $=\textbf{SAE}(h_{B}P_{\downarrow}+b_{\downarrow};\theta)P_{\uparrow}+b_{\uparrow}$
$=\Big{(}\sigma([h_{B}P_{\downarrow}+b_{\downarrow}]W_{e}+b_{e})W_{d}+b_{d}\Big{)}P_{\uparrow}+b_{\uparrow}$
$=\sigma(h_{B}[P_{\downarrow}W_{e}]+[b_{\downarrow}W_{e}+b_{e}])[W_{d}P_{\uparrow}]+[b_{d}P_{\uparrow}+b_{\uparrow}]$
$=\textbf{SAE}(h_{B};\theta^{\prime}),$
@@
$\|\textbf{SAE}(h_{B};\theta^{\prime})-h_{B}\|_{2}$ (13)
@@ B.2 Probes and Steering Vectors
To transfer a linear probe from $A$ to $B$, we simply apply the probe on the down-stitched residual stream $\mathcal{T}_{\downarrow}(B^{\ell_{B}}(t))$. If the probe is linear and has normal direction $w$, it is equivalent to using the vector $wP_{\downarrow}^{T}$ in $B$. Steering vectors are just transferred directly using $P_{\uparrow}$ without the bias.